OpenStudy (anonymous):

[6.01] The graph of the following system of equations is −2x + y = 3 −4x + 2y = 6 Overlapping lines Parallel lines Intersecting lines

OpenStudy (mertsj):

What is the slope of the first line?

OpenStudy (anonymous):

3

OpenStudy (mertsj):

The slope is -b/2a so the slope of the first line is m when you have the equation in y=mx+b form

OpenStudy (mertsj):

Do you see that if you multiply the first equation by 2 you will get the second equation?

OpenStudy (anonymous):

I see but how do i set up the second equation?

OpenStudy (mertsj):

You don't. You now know that the two equations are the same and therefore the two lines are the same line.

OpenStudy (anonymous):

so they make it Overlapping lines right??????

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

thank you so much. can you help me one more time if you like

OpenStudy (mertsj):

With what?

OpenStudy (anonymous):

with this [6.03] Solve the following system. x − y = 10 x + y = 8 (−9, 1) (9, −1) (9, 1) (−1, 9) i been taking to work it out but i think the answer is 9 -1 can you help me??

jimthompson5910 (jim_thompson5910):

(9,-1) is correct Check: x - y = 10 9 - (-1) = 10 .. replace x with 9, y with -1 9 + 1 = 10 10 = 10 ... works ----------- x + y = 8 9 + (-1) = 8 .. replace x with 9, y with -1 9 - 1 = 8 8 = 8 ... works both equations are true when x = 9 and y = -1, so (9,-1) is confirmed to be the solution

OpenStudy (anonymous):

I got that right lol

jimthompson5910 (jim_thompson5910):

nice job

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (anonymous):

[6.03] If you wanted to eliminate the y variable by adding the two equations in the following system, what could be your first step? 3x + y = 6 4x − 5y = 10 Multiply the first equation by 5. Multiply the first equation by −5. Multiply the first equation by −4. Add the two equations together.

jimthompson5910 (jim_thompson5910):

if you just added the equations as they are, then the y terms would add to this y + (-5y) = y - 5y = -4y

jimthompson5910 (jim_thompson5910):

but you want to make it 0y so the y terms go away when you add them

OpenStudy (anonymous):

so be this Add the two equations together.

jimthompson5910 (jim_thompson5910):

that doesn't eliminate the y variable, you have to do something before you add the equations

jimthompson5910 (jim_thompson5910):

y and -5y do not add to 0y

OpenStudy (anonymous):

Multiply them first

jimthompson5910 (jim_thompson5910):

by what

OpenStudy (anonymous):

-5y to both said

jimthompson5910 (jim_thompson5910):

how do you go from y to 5y

OpenStudy (anonymous):

idk i told you i suck at math

jimthompson5910 (jim_thompson5910):

well that's why practice is important, you'll get better the more you practice

jimthompson5910 (jim_thompson5910):

if we had y and -2y we can multiply the first 'y' by 2 to get 2y giving us: 2y and -2y they add to 2y + (-2y) = 2y - 2y = 0y

jimthompson5910 (jim_thompson5910):

see how I'm doing this?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

so what's the final answer?

OpenStudy (anonymous):

a

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

Thanks

jimthompson5910 (jim_thompson5910):

doing that will turn y into 5y then adding the two gives: 5y + (-5y) = 0y

jimthompson5910 (jim_thompson5910):

you're welcome